5X-7Y=1
2X+5Y=16
could you show the work as well please so i can learn how to do it?

Answer by Alan3354(69443)

5X-7Y=1
2X+5Y=16
--------
There's more than one way to do these.
I'll use elimination and substitution
------
5X-7Y=1
2X+5Y=16
Multiply each eqn to get the same coefficient for x or y:
10x-14y = 2 Eqn 1 times 2
10x+25y = 80 Eqn 2 times 5
------------- subtract
0x - 39y = -78
y = 2
------
Then sub for y into either eqn
2X+5Y=16
2x + 10 = 16
2x = 6
x = 3